what are the length and width of a rectangle if the length exceeds the width by 49 inches and the perimeter is 134 inches
Answer by fcabanski(1391)

For a rectangle Perimeter = P = 2L + 2W


In this problem call width = W and call length = W+49.


P = 134 = 2(W+49) + 2W = 2W + 98 + 2W


134 = 4W + 98


Subtract 98 from both sides.


36 = 4W


Divide both sides by 4.


9 = W


L = W + 49 = 9+49 = 58

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com